OpenStudy (anonymous):

can you explain how you got 12 hours for Carrie and 9 hours for Janice

OpenStudy (anonymous):

What is the context for this question?

OpenStudy (anonymous):

Carrie leave for vacation driving at a steady rate of 30 miles per hour. her friend Janice leaves from the same house to the same destination 3 hours later. she drives a rate of 40 miles per hour. how long after Janice leaves home will she catch up to Carrie

OpenStudy (anonymous):

Was there someone else that answered this previously? Because in your question you used the word "you" which makes me think there was someone else who answered it before and we could tag them in this?

OpenStudy (anonymous):

I mean I know how to get the answer but you're asking how they got it.

OpenStudy (anonymous):

yes but I want know how they got 12 hours for Carrie and 9 hours Janice that would be awesome if you can help me get the answer too.

OpenStudy (anonymous):

but I need help to get the answer for that math question

OpenStudy (mathlegend):

I would create a Table...

OpenStudy (anonymous):

ok why a table? Its an equation that I have to do

OpenStudy (mathlegend):

I see it best by creating a visual.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

When they meet up, the times are equal t = d/30= d/40 + 3 d/30 - d/40 =3 (40d - 30d)/1200 = 3 10 d = 3600 d = 360 d/30 = 12 Carry d/40 + 3 = 12 d/40 = 9 Janice

OpenStudy (mathlegend):